# Searching a Lightweight Network Architecture for Thermal Infrared Pedestrian Tracking

### v1

@@ Abstract
Manually-designed network architectures for thermal infrared pedestrian tracking (TIR-PT) require substantial effort from human experts. Neural networks with ResNet backbones are popular for TIR-PT. However, TIR-PT is a tracking task and more challenging than classification and detection. This paper makes an early attempt to search an optimal network architecture for TIR-PT automatically, employing single-bottom and dual-bottom cells as basic search units and incorporating eight operation candidates within the search space. To expedite the search process, a random channel selection strategy is employed prior to assessing operation candidates. Classification, batch hard triplet, and center loss are jointly used to retrain the searched architecture. The outcome is a high-performance network architecture that is both parameter- and computation-efficient. Extensive experiments proved the effectiveness of the automated method.
Index Terms:
@@
Thermal infrared, pedestrian tracking, neural architecture search.
I Introduction
@@
Thermal infrared pedestrian tracking (TIR-PT) is an active research field in computer vision and draws increasing interest in autonomous driving and underwater vehicles. It is generally treated as an object detection problem that intends to follow a specific pedestrian in TIR video sequences [1].
@@
Deep convolutional neural networks (CNNs) have been widely used in TIR-PT to extract features. Most of the CNNs of TIR-PT adopt typical networks such as AlexNet [2], VGGNet [3], and ResNet [4] as the backbone architecture, then integrate more layers to use deep features. For example, MMNet [5] implemented a multi-task matching network based on the AlexNet to learn object-specific discriminative and fine-grained correlation feature maps of TIR pedestrians. LMSCO [6] integrate motion information of the pedestrian into VGGNet to overcome the background clutter and motion blur of the TIR image. ASTMT [7] designed an aligned spatial-temporal memory network based on ResNet to take advantage of learning scene information for pedestrian localization. The well-performance handcrafted neural networks require substantial effort from human experts. However, these backbones were initially designed for image classification and thus have many redundant parts when transferred to different tasks. Since TIR-PT is an instance-level detection task, we hope to find a specific architecture suitable for it.
@@
Neural architecture search (NAS) has recently attracted wide attention in computer vision [8, 9]. NAS is an important advance that automates neural network designing. Given a search space, NAS methods aim to search a high-performance architecture with specific search strategies. In this paper, we attempt to search for an efficient network architecture for the TIR-PT task automatically. Conventional NAS methods cost intensive computation and memory. For instance, the reinforcement learning (RL)-based NASNet proposed by Zoph et al.  [10] takes 2000 GPU days to search and evaluate neural networks. AmoebaNet [11] is a well-known regularized evolution method. It costs 3150 GPU days to find a suitable architecture. We employ the recent state-of-the-art tracker DiMP [12] as our baseline and use a differentiable searching manner inspired by [13] to learn preferable network architecture for TIR-PT tasks. We relax the discrete search space to be continuous to optimize the architecture by gradient descent, and design a two-stage (search and retrain) method for TIR-PT, as shown in Fig.1. In the first stage, operation candidates in the search space are used to discover the basic architecture cells (or blocks). Then, the overall network architecture is constructed by stacking cells. Similar to the attention mechanism, apart from updating normal operation (such as convolutions) parameters $\omega$, we also need to update architecture parameters $\alpha$, an attention matrix representing the importance of each operation. Since it is a bi-level optimization problem, $\alpha$ and $\omega$ can be thus updated alternately. Though the differentiable method is relatively faster than other NAS, it still costs a lot of storage. Therefore, we randomly select channels and feed them into operation candidates to reduce computation, as suggested by [14]. The efficiency of the search process can be improved in this way. As for the second stage, we chose the most critical operation to form the final network architecture, similar to network pruning. We use joint supervision of classification, batch hard triplet, and center loss to learn more discriminative features to retrain the searched network architecture. Our contributions can be summarized as follows.
@@
  • We randomly select some channels and feed them into operation candidates to accelerate the search process.
  • We explore joint supervision to retrain the searched architecture to handle the challenging scenarios of TIR-PT.
  • We conduct extensive experiments to demonstrate the effectiveness and efficiency of our proposed method.
II The proposed method
II-A Basic search units
II-A1 Single-bottom Cells
Figure 1: Overview of the proposed TIR-PT method.
Figure 2: Pipeline of the single-bottom cell.
Figure 3: Pipeline of the dual-bottom cell.
@@ II-A1 Single-bottom Cells
II-A2 Dual-bottom Cells
@@
II-B Operation candidates
We define 8 operation candidates (search space) for each blue edge in Figs.2 and 3: none, 3$\times$3 max pooling, 3$\times$3 average pooling, 3$\times$3 depth-wise separable convolution, 5$\times$5 depth-wise separable convolution, skip connection, 3$\times$3 dilated convolution, 5$\times$5 dilated convolution. The normal and reduction cells are stacked repeatedly to form the final architecture. We keep the most crucial edge in inference to simplify architectures and reduce computations. We use a differential architecture search method to convert all discrete operations into a continuous search space by a softmax function. In this way, the architecture search task is simplified to learn a set of continuous variables,
$\bar{o}^{(i,j)}(x)=\sum_{o\in\mathcal{O}}\frac{\exp(\alpha_{o}^{(i,j)})}{\sum_{o^{\prime}\in\mathcal{O}}\exp(\alpha_{o^{\prime}}^{(i,j)})}o(x)$ (1)
@@ II-B Operation candidates
The continuous variables $a_{o}^{(i,j)}$ indicates the mixing weight associated with the operation $o$ between nodes $i$ and $j$ and forms a matrix $\alpha$. For Instance, dual-bottom cells are encoded as a matrix with 14 rows (edges) and 8 columns (operation candidates), while single-bottom cells are encoded as a matrix with 10 rows and 8 columns. The matrix represents the topology structure for a cell, named architecture parameters. Other operation parameters are marked as $\omega$. As for $\alpha$, it needs to optimize two kinds of the matrix where $\alpha_{1}$ is shared by all the normal cells and $\alpha_{2}$ is shared by all the reduction cells. Searching architecture parameters $\alpha$ and the operation parameters $\omega$ is a bi-level optimization problem, they are thus updated alternately.
II-C Random Channel Selection
Figure 4: Illustration of the random channel selection.
@@ II-C Random Channel Selection
$\bar{o}^{(i,j)}(x;r^{(i,j)})$ $=\sum_{o\in\mathcal{O}}\frac{\exp(\alpha_{o}^{(i,j)})}{\sum_{o^{\prime}\in\mathcal{O}}\exp(\alpha_{o^{\prime}}^{(i,j)})}o(r^{(i,j)}x)$ (2)
$+(1-r^{(i,j)})x$
where $\alpha$ is the searched architecture. $r^{(i,j)}x$ and $(1-r^{(i,j)})x$ respectively represent selected channels and the rest channels.
II-D Joint Supervision
@@
$\mathcal{L}_{BHTri}=\sum_{m=1}^{M}\sum_{n=1}^{N}$ $\big{[}\delta+\max_{p=1,\ldots,N}\mathcal{D}(f_{\alpha}(x_{m}^{n}),f_{\alpha}(x_{m}^{p}))$ (4)
$-\min_{\begin{subarray}{c}s=1,\ldots,M\\ t=1,\ldots,N\\ s\neq m\end{subarray}}\mathcal{D}(f_{\alpha}(x_{m}^{n}),f_{\alpha}(x_{s}^{t}))\big{]}_{+}$
@@
$\mathcal{L}_{CT}=\sum_{u=1}^{\mathcal{B}}\mathcal{D}(f_{\alpha}(x_{h_{u}}),f_{\alpha}(x_{y_{u}}))$ (5)
@@
$\mathcal{L}=\mathcal{L}_{CLS}+\mathcal{L}_{BHTri}+\gamma\mathcal{L}_{CT}$ (6)
where $\gamma$ is used for balancing the center loss.
III Experiments
III-A Datasets and Metrics
III-A1 Datasets
@@ III-A2 Metrics
III-B Implementation Details
In the searching process, random erasing, normalization, random horizontal flipping with 0.5, and padding the resized image 10 pixels with zero values are applied as the data augmentation. We hold out half of the training data to update architecture parameters $\alpha$. The other half of the training data can be viewed as a search validation set used to update operation parameters $\omega$. We utilize Adam [17] as the optimizer for $\alpha$, with an initial learning rate of 0.02, decay rates $\beta_{1}=0.5$ and $\beta_{2}=0.999$. We use momentum Stochastic Gradient Descent (SGD) to optimize the weights $\omega$. The initial learning rate starts from 0.1 and anneals down to 0.001 following a cosine schedule without restart for each epoch [18]. We search architecture for 200 epochs.
In the retraining process, we apply horizontal flip and normalization as data augmentation. SGD with momentum is deployed as the optimizer with an initial learning rate of 0.1 decayed with a cosine annealing. Each batch contains 8 pedestrians ($M$), and each pedestrian has 8 image frames ($N$), resulting in a batch size of 64.
The training dataset of LSOTB-TIR [15] is utilized to search and retrain the network architecture, we resize each training and test regions to 128$\times$128 and 256$\times$256. Except for the aforementioned details, the rest of the hyperparameters and configurations are are the same as DiMP [12]. We implement our method in Python using PyTorch with an instance with an ${}^{\circledR}$ ${}^{\circledR}$ Gold 6148v4 CPU @ 2.2 GHz CPU with 256 GB RAM and a ${}^{\circledR}$ ${}^{\circledR}$ V100-SXM2 GPU with 16 GB VRAM.
Figure 5: Illustration of the searched normal single-bottom cell.
Figure 6: Illustration of the searched reduction single-bottom cell.
Figure 7: Illustration of the searched normal dual-bottom cell.
Figure 8: Illustration of the searched reduction dual-bottom cell.
TABLE I: Ablation studies of the proposed method on the LSOTB-TIR benchmark dataset. S-c, D-c, CH, $\mathcal{L}_{CLS}$, $\mathcal{L}_{BHTri}$, and $\mathcal{L}_{CT}$ denote the single-bottom cell, dual-bottom cell, random channel selection, classification loss, batch triplet loss, and center loss, respectively.
Method configurations Precision Success Norm. Precision Params FLOPs GPU days
S-c D-c CH $\mathcal{L}_{CLS}$ $\mathcal{L}_{BHTri}$ $\mathcal{L}_{CT}$
0.768 0.640 0.697 2.96 M 0.78 G 3.3

### v2

@@ Abstract
Manually-designed network architectures for thermal infrared pedestrian tracking (TIR-PT) require substantial effort from human experts. AlexNet and ResNet are widely used as backbone networks in TIR-PT applications. However, these architectures were originally designed for image classification and object detection tasks, which are less complex than the challenges presented by TIR-PT. This paper makes an early attempt to search an optimal network architecture for TIR-PT automatically, employing single-bottom and dual-bottom cells as basic search units and incorporating eight operation candidates within the search space. To expedite the search process, a random channel selection strategy is employed prior to assessing operation candidates. Classification, batch hard triplet, and center loss are jointly used to retrain the searched architecture. The outcome is a high-performance network architecture that is both parameter- and computation-efficient. Extensive experiments proved the effectiveness of the automated method.
@@
Keywords: Thermal infrared, pedestrian tracking, machine learning, neural network
1 Introduction
@@
Thermal infrared pedestrian tracking (TIR-PT) is an active research field in computer vision and draws increasing interest in autonomous driving and underwater vehicles. It is generally treated as an object detection problem that intends to follow a specific pedestrian in TIR video sequences [1, 2].
@@
Convolutional neural networks (CNNs) have been widely used in TIR-PT to extract features. Most of the CNNs of TIR-PT adopt typical networks such as AlexNet [3] and ResNet [4] as the backbone architecture, then integrate more layers to use deep features.For example, MMNet [5] implemented a multi-task matching network based on AlexNet to learn object-specific discriminative and fine-grained correlation feature maps of TIR pedestrians. LMSCO [6] integrated the motion information of the pedestrian into VGGNet to overcome the background clutter and motion blur of the TIR image. ASTMT [7] designed an aligned spatial-temporal memory network based on ResNet to take advantage of learning scene information for pedestrian localization. The well-performance handcrafted neural networks require substantial effort from human experts. However, these backbones are initially designed for image classification and object detection and thus have many redundant parts when transferred to different tasks. Since TIR-PT is an instance-level detection task, we hope to find a specific architecture suitable for it.
Neural architecture search (NAS) has recently attracted wide attention in classification [8], detection [9], tracking [10], and segmentation [11]. It is an important advance that automates neural network designing. In this paper, we attempt to search for an efficient network architecture for the TIR-PT task automatically. Given a search space, NAS methods aim to search a high-performance architecture with specific search strategies. Conventional NAS methods cost intensive computation and memory. For instance, the reinforcement learning (RL)-based NASNet proposed by Zoph et al.  [12] takes 2000 GPU days to search and evaluate neural networks. AmoebaNet [13] is a well-known regularized evolution method. It costs 3150 GPU days to find a suitable architecture.
Figure 1: Overview of the proposed TIR-PT method.
@@
Inspired by the aforementioned works, in this paper, we employ the discriminative tracking model DiMP [14] as our baseline and use a differentiable searching manner inspired by [10] to learn a preferable network architecture for TIR-PT tasks. We relax the discrete search space to be continuous to optimize the architecture by gradient descent, and design a two-stage (search and retrain) method for TIR-PT, as shown in Fig.1. In the first stage, operation candidates in the search space are used to discover the basic architecture cells (or blocks). Then, the overall network architecture is constructed by stacking cells. Similar to the attention mechanism, apart from updating normal operation (such as convolutions) parameters $\omega$, we also need to update architecture parameters $\alpha$, an attention matrix representing the importance of each operation. Since it is a bi-level optimization problem, $\alpha$ and $\omega$ can be thus updated alternately. Though the differentiable method is relatively faster than other NAS, it still costs a lot of storage. Therefore, we randomly select channels and feed them into operation candidates to reduce computation, as suggested by [15]. The efficiency of the search process can be improved in this way. As for the second stage, we chose the most critical operation to form the final network architecture, similar to network pruning. We use joint supervision of classification, batch hard triplet, and center loss to learn more discriminative features to retrain the searched network architecture. Our contributions can be summarized as follows.
  • We made an early attempt to search automatically for an effective network architecture for the TIR-PT task. The architecture learned on the LSOTB-TIR dataset can be directly transferred to the PTB-TIR dataset with high performance.
@@
  • Instead of calculating all channels, we randomly select a part of the channels and feed them into operation candidates to accelerate the search process.
  • Considering the challenging scenarios of TIR-PT, we explore joint supervision of classification, batch hard triplet, and center loss to retrain the searched network architecture.
  • Extensive experimental results demonstrated the effectiveness and efficiency of our proposed method.
The remainder of this paper is organized as follows. Relevant TIR-PT and NAS methods are introduced in Section 2. Section 3 details our methodology for searching for an optimal neural network architecture for TIR-PT. In the following section, Section 4, experimental results and discussions demonstrate that the proposed method is efficient for TIR-PT. Finally, a brief conclusion is provided in Section 5.
2 Related works
2.1 Thermal infrared pedestrian tracking
TIR-PT represents a significant advancement in surveillance, security, and urban management technologies. Its ability to operate in low-light conditions, independence from weather constraints, and privacy-sensitive imaging make it a versatile tool in various applications. Early TIR trackers used handcrafted features [16], but the DSST-tir tracker showed that deep features are more effective [17]. Consequently, more TIR trackers started incorporating deep features. MCFTS [18], for instance, uses VGGNet layers for an ensemble tracker, while HSSNet [19] transfers a deep Siamese framework from a large video dataset to the TIR domain. Recent advancements include MLSSNet [20], which excels in distinguishing distractors, and MMNet [5], which innovates dual-level deep representation learning. Prevailing backbone networks utilized by existing TIR-PT methods have been designed manually by human experts, which is time-consuming and requires a lot of prior knowledge from experts. Our goal is to automatically search an architecture for the TIR-PT task.
2.2 Network architecture search
In recent years, NAS has attracted growing attention and plays a vital role in automatic machine learning (AutoML). The goal is to search for architectures automatically to replace conventional manual ones. Early approaches take a large amount of computation. Zoph et al. [21] use a recurrent neural network (RNN) to encode the network architecture, which takes 800 GPUs for 28 days, resulting in 22,400 GPU hours to search for a suitable architecture on the CIFAR-10 dataset. Further, NASNet, proposed by Zoph et al. [12], uses 500 GPUs across four days, resulting in 2,000 GPU hours. Metacontroller is implemented as an RL agent in [22]. It learns to take action for network transformation (widen layer, insert layer, add skip-connections, etc.) to explore the architecture space. ENAS, proposed by Pham et al. [23], speeds up NAS by sharing parameters among child models during the search procedure. Some methods try to reduce computation costs. In particular, Liu et al. [24] use a differentiable method to relax the discrete operations into continuous variables. The essential edges with high attention-weighted scores will be left during inference. Xu et al. [15] explore larger batch sizes for either speedup or higher stability with partial-connected mechanism.
Our method is closest to DARTS [24]. For RL-based and evolutionary-based methods, feedback (a.k.a., reward) is obtained after a long training. In contrast, feedback (a.k.a., loss) in the gradient-based method is given in every iteration, which is more instant. Therefore, the optimization of the gradient-based method is potentially more efficient. The discrete search space can be transferred to a continuous matrix with softmax, enabling the search to progress end-to-end.
3 Methodology
Figure 2: A single-bottom cell contains six nodes. Node S is the input feature map. Nodes 0, 1, 2, 3 are intermediate feature maps. The last output node is depth-wise concatenation of four intermediate nodes. Each blue edge denotes eight operation candidates.
Figure 3: A dual-bottom cell contains seven nodes. Nodes S0 and S1 are feature maps outputted from the two previous layers.
3.1 Basic search units
3.1.1 Single-bottom cells
@@
3.1.2 Dual-bottom cells
@@
3.2 Operation candidates
We define eight operation candidates (a.k.a., search space) for each blue edge in Figs.2 and 3: none, 3$\times$3 max pooling, 3$\times$3 average pooling, 3$\times$3 depth-wise separable convolution, 5$\times$5 depth-wise separable convolution, skip connection, 3$\times$3 dilated convolution, 5$\times$5 dilated convolution. The normal and reduction cells are stacked repeatedly to form the final architecture. We keep the most crucial edge in inference to simplify architectures and reduce computations. We use a differential architecture search method to convert all discrete operations into a continuous search space by a softmax function. In this way, the architecture search task is simplified to learn a set of continuous variables,
$\bar{o}^{(i,j)}(x)=\sum_{o\in\mathcal{O}}\frac{\exp(\alpha_{o}^{(i,j)})}{\sum_{o^{\prime}\in\mathcal{O}}\exp(\alpha_{o^{\prime}}^{(i,j)})}o(x),$ (1)
@@ 3.2 Operation candidates
Figure 4: Architecture parameters $\alpha$ for a dual-bottom cell. Rows and columns represent 14 edges and 8 operation candidates, respectively.
The continuous variables $a_{o}^{(i,j)}$ indicates the mixing weight associated with the operation $o$ between nodes $i$ and $j$ and forms a matrix $\alpha$. For Instance, dual-bottom cells are encoded as a matrix with 14 rows (edges) and 8 columns (operation candidates) as shown in Fig.4. The matrix represents the topology structure for a cell, named architecture parameters. Other operation parameters are marked as $\omega$. The network architecture samples with single- and dual-bottom cells are illustrated in Figs.5 and 6, respectively. Size of input image is $C\times H\times W$, specifically $3\times 256\times 256$ and $3\times 128\times 128$ for training and test regions. Blue blocks represent reduction cells, which reduce the size of feature maps and double the number of channels. Green blocks are normal cells utilized to enhance the expressive performance of the network. Searching architecture parameters $\alpha$ and the operation parameters $\omega$ is a bi-level optimization problem, they are thus updated alternately. As for $\alpha$, it needs to optimize two kinds of the matrix where $\alpha_{1}$ is shared by all the normal cells and $\alpha_{2}$ is shared by all the reduction cells.
Figure 5: The overview of network architecture sample with only one input (a.k.a., single-bottom).
Figure 6: The overview of network architecture sample with two inputs (a.k.a., dual-bottom).
3.3 Random channel selection
@@
$\bar{o}^{(i,j)}(x;r^{(i,j)})=\sum_{o\in\mathcal{O}}\frac{\exp(\alpha_{o}^{(i,j)})}{\sum_{o^{\prime}\in\mathcal{O}}\exp(\alpha_{o^{\prime}}^{(i,j)})}o(r^{(i,j)}x)+(1-r^{(i,j)})x,$ (2)
where $\alpha$ is the searched network architecture, $r^{(i,j)}x$ and $(1-r^{(i,j)})x$ respectively represent selected channels and the rest channels.
Figure 7: Illustration of the random channel selection. Instead of calculating all channels, we select a part of channels and feed them into operation candidates, then concatenate with the rest of the unchanged channels.
3.4 Joint supervision
@@
$\mathcal{L}_{BHTri}=\sum_{m=1}^{M}\sum_{n=1}^{N}\big{[}\delta+\max_{p=1,\ldots,N}\mathcal{D}(f_{\alpha}(x_{m}^{n}),f_{\alpha}(x_{m}^{p}))-\min_{\begin{subarray}{c}s=1,\ldots,M\\ t=1,\ldots,N\\ s\neq m\end{subarray}}\mathcal{D}(f_{\alpha}(x_{m}^{n}),f_{\alpha}(x_{s}^{t}))\big{]}_{+},$ (4)
@@
$\mathcal{L}_{CT}=\sum_{u=1}^{\mathcal{B}}\mathcal{D}(f_{\alpha}(x_{h_{u}}),f_{\alpha}(x_{y_{u}})),$ (5)
@@
$\mathcal{L}=\mathcal{L}_{CLS}+\mathcal{L}_{BHTri}+\gamma\mathcal{L}_{CT},$ (6)
where $\gamma$ is used for balancing the center loss and set to 0.5.
4 Experiments
Figure 8: Illustration of the searched normal single-bottom cell. In order to simplify architecture and reduce calculations, each intermediate node takes the most important edge as input. It is similar to attention-based pruning.
Figure 9: Illustration of the searched reduction single-bottom cell. The function of reduction cell is to reduce the size of feature maps and double channels. Intermediate layer only takes the most important edges as input. The last output is depth-wise concatenation of four intermediate nodes.
4.1 Datasets and metrics
@@
4.2 Implementation details
In the searching process, random erasing, normalization, random horizontal flipping with 0.5, and padding the resized image 10 pixels with zero values are applied as the data augmentation. We hold out half of the training data to update architecture parameters $\alpha$. The other half of the training data can be viewed as a search validation set used to update operation parameters $\omega$. We utilize Adam [27] as the optimizer for $\alpha$, with an initial learning rate of 0.02, decay rates $\beta_{1}=0.5$ and $\beta_{2}=0.999$. We use stochastic gradient descent (SGD) with momentum to optimize the weights $\omega$. The initial learning rate starts from 0.1 and anneals down to 0.001 following a cosine schedule without restart for each epoch [28]. We search architecture for 200 epochs. In the retraining process, we apply horizontal flip and normalization as data augmentation. SGD with momentum is deployed as the optimizer with an initial learning rate of 0.1 decayed with a cosine annealing. Each batch contains 8 pedestrians (a.k.a.$M$), and each pedestrian has 8 image frames (a.k.a.$N$), resulting in a batch size of 64.
The training dataset of LSOTB-TIR [25] is utilized to search and retrain the network architecture, we resize each training and test regions to 128$\times$128 and 256$\times$256. Except for the aforementioned details, the rest of the hyperparameters and configurations are are the same as DiMP [14]. We implement our method in Python using PyTorch with a cloud server with an Intel® Xeon® Gold 6148v4 CPU @ 2.2 GHz CPU with 256 GB RAM and a NVIDIA® Tesla® V100-SXM2 GPU with 16 GB VRAM.
4.3 Ablation studies
Figure 10: Illustration of the searched normal dual-bottom cell. To balance efficiency and parameters, each intermediate layer has two important inputs according architecture matrix.
Figure 11: Illustration of the searched reduction dual-bottom cell. It takes two important layers as current input according to architecture parameters.
Figure 12: Illustration of the final searched network architecture. “R” and “N” indicate the searched reduction and normal dual-bottom cells, respectively.
Table 1: Ablation studies of the proposed method on the LSOTB-TIR benchmark dataset. #Dims and #Cells represent the dimensionality of the output feature map and the number of stacked cells of the network architecture, respectively. Single, Dual, Channel, $\mathcal{L}_{CLS}$, $\mathcal{L}_{BHTri}$, and $\mathcal{L}_{CT}$ denote the single-bottom cell, dual-bottom cell, random channel selection, classification loss, batch triplet loss, and center loss, respectively. The best two results are shown in red and blue fonts.
#Cells #Dims Configurations Precision $\uparrow$ Success $\uparrow$ Norm. Precision $\uparrow$ Params $\downarrow$ FLOPs $\downarrow$ GPU days $\downarrow$
Single Dual Channel $\mathcal{L}_{CLS}$ $\mathcal{L}_{BHTri}$ $\mathcal{L}_{CT}$
16 512 0.768 0.640 0.697 2.96 M 0.78 G 3.3